OpenStudy (samigupta8):

force acting on a particle in conservative force field is 1. F=(2i+3j) 2.F=(2x i+ 2y j) 3.F=(yi+xj) find the potential energy function if it is zero at origin

OpenStudy (anonymous):

\[E=-\int\limits_{0}^{x}F_xdx-\int\limits_{0}^{y}F_ydy\]

OpenStudy (samigupta8):

\[ E=-\int\limits_{0}^{r}Fdr=-\int\limits_{0}^{r}(2i+3j)dr\]

OpenStudy (anonymous):

dr=dxi+dyj+dzk

OpenStudy (anonymous):

But Fdr is a dot product, then it is Fx·dx+Fy·dy

OpenStudy (samigupta8):

yyaa ryt then how are u able to solve it furthermore u cnnot put the value of i and j in integration can u?

OpenStudy (anonymous):

F=2x+3y

OpenStudy (anonymous):

of course minus F

OpenStudy (anonymous):

You have to integrate: \[(F_x \hat i+F_y \hat j)·(\hat i dx+\hat j dy)=F_xdx+F_ydy\]

OpenStudy (samigupta8):

kkk.....

OpenStudy (anonymous):

Case number 2 \[E=-\int\limits_{0}^{r}(2x \hat i+2y \hat j)(\hat i dx+\hat j dy)=-\int\limits_{0}^{x}2xdx-\int\limits_{0}^{y}2ydy=-(x^2+y^2)\]

OpenStudy (samigupta8):

what will be dr in 1st case

OpenStudy (anonymous):

dr=dxi+dyj+dzk

OpenStudy (anonymous):

dr is always ^idx + ^jdy what changes is the force Fxi^+Fyj^

OpenStudy (anonymous):

In first case Fx=2 and Fy=3, the you have to integrate -2dx-3dy---->E=-(2x+3y)

OpenStudy (samigupta8):

how will we solve for third part

OpenStudy (anonymous):

integrate -(Fxdx+Fydy)= integrate -ydx-xdy----->E=-(yx+xy)=-2yx

OpenStudy (samigupta8):

bt in answer it is given as only -xy

OpenStudy (anonymous):

Then they have given the wrong answer

OpenStudy (samigupta8):

yaa u mst be right bcuz i m also getting the same answer aftr its integration